Erin has red, blue, and yellow barrettes. She has 5 times as many blue barrettes as red barrettes, and 2 times as many yellow barrettes as red barrettes. Altogether, she has 24 barrettes. How many of each color does she have?
Answer by gonzo(654)

let b = blue
let r = red
let y = yellow
-----
b + r + y = 24
b = 5r
y = 24
substitute 5r for b and 2r for yellow:
5r + r + 2r = 24
8r = 24
r = 3
b = 5r = 15
y = 2r = 6
b + r + y = 24
15 + 3 + 6 = 24
24 = 24 - true
answer is:
3 red
15 blue
6 yellow
-----